[35844] in Hotline Meeting

home help back first fref pref prev next nref lref last post

Requested Shutdown - All of E15 - Filters Changed - 11/19/96

daemon@ATHENA.MIT.EDU (Dennis J. Irving)
Mon Nov 18 09:19:42 1996

Date: Mon, 18 Nov 1996 09:20:59 -0500
To: OPCTEMP2%MITPPL.bitnet@mitvma.mit.edu
From: dirving@media.mit.edu (Dennis J. Irving)
Cc: bldg-help@media.mit.edu, shutdowns@MIT.EDU, bv@media.mit.edu,
        haller@media.mit.edu, jroll@MIT.EDU

Andrea,

The air handler filter change mentioned below is ok with the Facilities
Services dept. of the Media Lab.

The following requested shutdown for E15 for 11/19 is for your
>information and comment. Would you please let me know if this has your
>approval or disapproval.
>
>Thanks very much.
>
>Andrea Nichols
>Operations Center
>8-7872
>
>Building:       E15 - All Floors, Film Video Theater, Lobby and Atrium
>System:         Air Handler 1, 2, 3 and 4, Supply Fans 1A and 1B
>Date:           11/19/96
>Day:            Tuesday
>Time:           9:00 AM - 10:15 AM
>Duration:       1-1/4 hours
>Reason:         Change Filters
>Effect:         Loss of supply-ventilation as follows: for AH1 A & B -
>                these will be down for 20 minutes each. For AH 2 and
>                3, these will be down for 15 minutes each.
>
>                Note: While one air handler is down, all others will be
>                operational.
>

Dennis Irving
Facilities Coordinator
253-0684, E15-463F



home help back first fref pref prev next nref lref last post